jsinrod:

What is the result when the number 16 is increased by 75%?

jsinrod:

...

iYuko:

Start with finding what 75% of 16 is by doing 16 x .75

cutebunny123:

The result when the number 16 is increased by 75% is 12. This is how I got my answer 12 well, 75% in 16 is 12. The, I multpied 75x100 with 16 (75x100)*16 = (75*16)/100 = 12. I hope this helps you!

iYuko:

@cutebunny123 wrote:
The result when the number 16 is increased by 75% is 12. This is how I got my answer 12 well, 75% in 16 is 12. The, I multpied 75x100 with 16 (75x100)*16 = (75*16)/100 = 12. I hope this helps you!
You need to add 12 to 16 and get \[28\]